8. How many haploid sets of chromosomes are present in an individual cell that is pentaploid

5n

A) 2

B) 3

C) 4

D) 5

E) It is impossible to tell with the information given.

9. The process that leads to development haploid gamete is best described as ________.

A) segregation

B) independent assortment

C) Mendelian inheritance

D) replication

E) dominance or recessiveness

10. If an F2 generation from a self-cross always yields offspring in a 3:1 phenotypic ratio, which of the

following P crosses could have occurred?

A) AA × aa

B) AA × AA

C) aa × aa

D) Aa × AA

E) aa × Aa

11. In a mating between individuals with the genotypes

I

Ai

×

ii

, what percentage of the offspring are

expected to have the O blood type?

A) 0%

B) 25%

C) 50%

D) 75%

E) 100%

12. Which of the following is

not

consistent with Erwin Chargaff’s findings?

a. (A + G) = (C + T)

b. (A + C) = (G + T)

c. A = T

d. (C + G) = (A + T)

e. (A + G)/(C + T)

1

13. The virus phi X 174 infects

Escherichia coli

.

Its base composition is as follows:

A = 24.0 %

G = 23.3 %

C = 21.4 %

T = 31.3 %

What conclusions can you draw about phi X 174 ?

a.

It has a double-stranded D N A genome.

b. It has a single-stranded D N A genome.

c.

It has a double-stranded R N A genome.

d. It has a single-stranded R N A genome.

e. A must bind to G and C to T in this virus.
Answer #1

Answer:

8). D) 5

Explanation:

Penta = 5

Pentaploid means that the presence of 5 (haploid) sets of chromosomes.

9). B). Independent assortment

10). A). AA x aa

Explanation:

AA x aa ---Parents

Aa (round)---F1

Aa x Aa ---F1 selfcross

A

a

A

AA

Aa

a

Aa

aa

Dominant trait = AA+Aa = ¾

Recessive trait = aa = ¼

Phenotypic ratio = 3:1

11). C). 50%

Explanation:

IAi (A type) x ii (O type)---Parents

i

IA

IAi (A type)

i

ii (O type)

A type = 50%

O type = 50%

12). d. (C + G) = (A + T)

13). b. It has a single-stranded D N A genome.

Explanation:

Thymine is present; hence, it is DNA. The percentages of nitrogen bases do not follow Chargaff’s rule; hence, it is not double stranded.